How many moles are there in 8.25 x 1026 molecules of methane, CH4?

1
Expert's answer
2022-05-18T02:19:03-0400

Solution:

One mole of any substance contains 6.02×1023 atoms/molecules.

Therefore,

Moles of CH4 = (8.25×1026 molecules) × (1 mol / 6.02×1023 molecules) = 1370 mol = 1.37×103 mol

Moles of CH4 = 1.37×103 mol


Answer: 1.37×103 moles of methane (CH4)